OpenStudy (anonymous):

Evaluate: 5(6-3)-1x6

mathslover (mathslover):

\[\Huge{\mathbb{Welcome}\textbf{To }\mathbb{Open}\textbf{Study}}\] what is 6-3 ?

OpenStudy (lgbasallote):

so what's your answer here?

OpenStudy (anonymous):

3

mathslover (mathslover):

ok so 5(6-3)=5(3) right ?

OpenStudy (anonymous):

i came up with 84 when i worked it on paper and 9 when i worked it on my calculator.

mathslover (mathslover):

Ok so 5*3 = ?

OpenStudy (anonymous):

15

mathslover (mathslover):

1*6 = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

15-6x

mathslover (mathslover):

right now what is 15-6 ?

OpenStudy (anonymous):

9

mathslover (mathslover):

hence 9 is your answer congrats

mathslover (mathslover):

Check it again that why were u getting 84 instead of 9

OpenStudy (anonymous):

i was subtracting to early. Thank you.